Circular No. 6147

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

X-RAY NOVA 1994 IN SCORPIUS B. A. Harmon, S. N. Zhang, W. S. Paciesas, and G. J. Fishman report

for the Compton Observatory BATSE Team: “Since the bright outburst of the x-ray transient GRO J1655–40 (∼1 Crab, band 20–100 keV) in hard x-rays in 1994 Dec. (ending ∼ Dec. 23), the source has exhibited two additional episodes of variable high-energy emission. The first began around 1995 Jan. 12 (IAUC 6128) and ended ∼ Jan. 25. The second began around Feb. 12 and is still in progress. Both episodes show daily variations from < 50 mCrab to > 300 mCrab (band 20–100 keV). As in the 1994 outbursts, the spectrum is a power law extending to at least 200 keV. However, the best-fit spectral indices are harder, ranging between −2.5 and −1.9. No unusual hard-x-ray activity was observed in the BATSE large-area detectors during the Mir/Kvant/TTM measurements on Feb. 22 and 24 (IAUC 6143; R. Sunyaev, private communication).”

PSR 1951+32 P. V. Ramanamurthy, D. L. Bertsch, B. L. Dingus, J. A. Esposito,

C. E. Fichtel, S. D. Hunter, C. v. Montigny, R. Mukherjee, and D. J. Thompson, Goddard Space Flight Center, NASA; J. M. Fierro, Y. C. Lin, P. F. Michelson, and P. L. Nolan, Stanford University; G. Kanbach, H. A. Mayer-Hasselwander, and M. Merck, Max-Planck-Institut für Extraterrestrische Physik; D. A. Kniffen, Hampden-Sydney College; A. G. Lyne, University of Manchester; and J. R. Mattox, University of Maryland, report: “We detected a sixth high-energy, γ-ray pulsar, PSR 1951+32, pulsating in γ-rays at energy > 100 MeV with the same 39.5-ms periodicity as detected at radio wavelengths, using the data obtained during May 1991– July 1994 by the Energetic Gamma Ray Experiment Telescope aboard the Compton Gamma Ray Observatory. Though seen only as a weak source amidst the high background of galactic disk emission, the pulsation in highenergy γ-rays is clearly seen, with the null (no pulsed emission) hypothesis being probable only at a level of < 10^{−9}. The pulsed radiation has a photon spectral index of −(1.74±0.11), and the phase-averaged integral flux of pulsed γ-rays is (1.6 ± 0.2) × 10^{−7} cm^{−2} s^{−1} at energy > 100 MeV. There is no evidence as yet for unpulsed emission from the object. The pulsar appears to have an efficiency of ∼ 0.0038 for converting its rotational energy loss into γ-rays at energy > 100 MeV.”

1995 March 8 Daniel W. E. Green